# PROFIT: A Specialized Optimizer for Deep Fine Tuning

### v2

@@ 3.3 Theoretical Considerations
We now list a few simple theoretical properties of PROFIT, along with sketches of proofs. We first prove that PROFIT is “correct”, in the sense that it is able to decrease the loss on the old model task/setting. We also add a discussion of when PROFIT might fail by enumerating all stable points of PROFIT, but argue why this is generally not an issue in practical settings.
{theorem}
@@
[] (Correctness on old data.) Take a model $\mathcal{M}(\mathbf{x};\theta)$ converged on data $\mathcal{X}_{\text{old}}$ with loss $L_{\text{old}}$, which we would now like to fine-tune on data $\mathcal{X}_{\text{new}}$ with loss $L_{\text{new}}$. Suppose that $\mathbf{O}^{(\text{ref})}$ takes $\theta\mapsto\theta^{\prime}$. A single step of PROFIT on batch $\mathbf{x}_{\text{new}}$ with a sufficiently small learning rates $\lambda_{\text{main}}$, $\lambda_{\text{ref}}$ will decrease $L_{\text{old}}(\mathcal{X}_{\text{old}})$ from its value at $\theta^{\prime}$.
Proof Sketch: If $\mathcal{M}$ is converged on data $\mathcal{X}_{\text{old}}$, then it must be at a local minimum of the old data. If the learning rate is sufficiently small, then $\nabla_{\theta^{\prime}}L_{\text{old}}$ must point in the $\Delta$ (as defined in Alg 1) direction, and so an update with the gradient $\Delta+\langle\Delta,\mathbf{g}\rangle$ will necessarily decrease $L_{\text{old}}$.∎
@@
The prior theorem establishes that PROFIT accomplishes precisely what it seeks to do: even if the old data is no longer available, PROFIT allows us to train as if we can still compute the full loss function of the old data. As the system moves further away from the old equilibrium, we are able to restore some of the function of that equilibrium through these regularized updates. Notably, there is a case where PROFIT leads to a trivial update.
{theorem}
@@
[] (Stable points.) Suppose a model has weights $\theta$ and $\mathbf{O}^{(\text{ref})}$ maps $\theta\mapsto\theta^{\prime}$ and $\mathbf{O}^{(ref)}$ is SGD. If we are not at a critical point of $L_{\text{new}}$, PROFIT will result in zero change in the model weights $\mathbf{w}$ if and only if ^$\hat{\nabla}_{\theta}L_{\text{new}}=\hat{\nabla}_{\theta^{\prime}}L_{\text{new}}$, where $\hat{\nabla}$ refers to the unit vector corresponding to $\nabla$.
@@
Proof: $\Delta$ in this case would point in the direction $-\nabla_{\theta}L_{\text{new}}$, at which point $\nabla_{\theta}L_{\text{new}}\perp\Delta=0$, and the total update from PROFIT will be just a restoration to $\theta$. If the gradient equality does not hold, $\nabla_{\theta}L_{\text{new}}\perp\Delta\neq 0$ will lead to a non-zero total update. ∎
{corollary}
@@
[] (Linearity forces a stable point.) In the situation defined by Theorem 3.3, PROFIT will encounter a stable point if the loss surface is perfectly linear between $\theta$ and $\theta^{\prime}$. ∎
@@
The prior theorem and corollary demonstrate that PROFIT will fail to move the system when the loss surface becomes exactly locally linear at a point. This almost never occurs for high-dimensional loss surfaces that exist for deep models, and even in the 2-dimensional toy example described in Section 4.1 we see a reasonable performance.
3.4 Assumptions and Limitations
@@
Orthogonalizing between $\mathbf{g}$ and $\Delta$ implies that we are interested in resolving the conflict between the original and fine-tuning datasets. If the original and fine-tuning distributions are significantly different, the orthogonalized gradient direction may not benefit either tasks and lead to destructive interference. Therefore, our interpretation holds only if the original and fine-tuning datasets are from proximal distributions. Assumption 1 may differ from typical fine-tuning literature and appear greatly limiting, but is exceedingly common in practice. For instance, a self-driving car may need to train a classification head for a new type of scooter, while also maintaining good performance on previously observed vehicles.
Assumption 2 is required since our method is exclusively designed for fine-tuning (which is a guiding principle). Nonetheless, we include it here to emphasize to the reader that PROFIT cannot be used for general model optimization.
As discussed in Sec 3.3, PROFIT requires inference on $n_{\text{ref}}+1$ batches (two for $n_{\text{ref}}=1$) per iteration, hence is slower than standard optimizers for fine-tuning. In practice, we observe that PROFIT converges earlier than corresponding baselines (which are allowed to train longer for fair comparison), which may mitigate this. We also note that typical fine-tuning settings are shorter and not as involved as training from scratch, which may further alleviate this concern. Another limitation is that our method would consume slightly more memory as a result of instantiating two optimizers ($\mathbf{O}^{\text{(ref)}}$ and $\mathbf{O}$).
@@ 4 Experiments
We now detail a number of experiments for PROFIT in diverse settings: low-dimensional toy setting to image classification, visual task adaptation, and large-scale motion prediction. We primarily focus on comparisons to standard fine-tuning with standard optimizers (on either the full model or just the model head), as those are - by a large margin - still the most commonly used fine-tuning methods in the industry due to their known performance and ease of implementation. We will show that PROFIT, while being easy to implement, provides a significant performance boost in all cases.
@@ 4.1 A Simple Toy Example
We present results in this section on a simple, low-dimensional toy example. We find that testing new algorithms in low-dimensional settings is always instructive and provides us with insight that cannot be gleaned from the noise of higher-dimensional experiments.
Namely, we fit a number of simple MLP models onto a 2D function. We pick the 2D function $f(\mathbf{x})=\sin(10|\mathbf{x}|)$, because both radially symmetric and periodic functions are generally more challenging for neural networks to fit. We further add $\mathcal{N}(0,1)$ noise to the output. Even though we want to fit to a low-dimensional example, it is still important for the problem to be difficult enough to see interesting behavior within our models, especially given the dimensionality requirements as discussed in Section 3.3. The original dataset consists of input-output pairs where input coordinates are drawn independently from $\mathcal{U}[-1.0,1.0]$, while the new dataset we wish to fine-tune on has input coordinates drawn from $\mathcal{U}[0.8,1.5]$. This is clearly also challenging because, although the domains of the two datasets overlap in the interval $[0.8,1.0]$, they are largely nonoverlapping.
Our MLP consists of three layers with weights of shape $[2,500]\rightarrow[500,500]\rightarrow[500,1]$, and we use RMSProp as our baseline optimizer. The baseline model is trained on the original data split only for 10000 steps at a learning rate of 1e-2, with fine-tuning runs trained at a learning rate of 5e-4 for 1500 steps. For exact details on the training procedure, please refer to the supplementary material.
Figure 2: Fine-tuning toy example visualizations. The top row shows ground truth distributions for the original, new, and combined datasets, while the bottom row depicts model predictions for different training strategies: training on original data only, head-only fine-tuning, full model fine-tuning, and using the proposed PROFIT optimizer for full model fine-tuning. This highlights the PROFIT optimizer’s effectiveness in retaining old task knowledge while adapting to new tasks. Best viewed in color.
Method Original Data Error ()($\downarrow$)( ↓ ) New Data Error ()($\downarrow$)( ↓ )
@@
Table 1: Results on a 2D fitting problem. A baseline is trained on just the original data domain and then fine-tuned at a lower learning rate on a new data domain (both the full model and the head only). This is compared to PROFIT, which is trained on the full model. As expected, there is a clear tradeoff between training on the full model versus the head only; the latter provides some protection against performance regression on the original data, but also is less able to adapt to the new data. PROFIT outperforms both baselines and shows impressive resilience in maintaining performance on the original split. All results have standard error $\leq 0.01$.
The results are visualized in Figure 2. Visually, the benefits of PROFIT are fairly pronounced; fine-tuning on the full model creates a warped shape bearing minimal resemblance to the original ground truth, and even fine-tuning on the head layer only still produces sizable shifts in the overall height of the output shape. PROFIT effectively remembers the original shape with only minimal regressions along the steep edges of the distribution.

### v3

@@ 3.2 Theoretical Considerations
We present key theoretical properties of PROFIT with proof sketches. First, we show that PROFIT is “correct” by decreasing the loss on the old task/setting. We also discuss potential failure cases by identifying all stable points of PROFIT, arguing that these are not problematic in practice.
Theorem 3.1.
@@
(Correctness on old data) Take a model $\mathcal{M}(\mathbf{x};\theta)$ converged on data $\mathcal{X}_{\text{old}}$ with loss $L_{\text{old}}$, which we would now like to fine-tune on data $\mathcal{X}_{\text{new}}$ with loss $L_{\text{new}}$. Suppose that $\mathbf{O}^{(\text{ref})}$ takes $\theta\mapsto\theta^{\prime}$. A single step of PROFIT in batch $\mathbf{x}_{\text{new}}$ with sufficiently low learning rates $\lambda_{\text{main}}$, $\lambda_{\text{ref}}$ will decrease $L_{\text{old}}(\mathcal{X}_{\text{old}})$ from its value at $\theta^{\prime}$.
Proof.
At a convergence point local minimum $\mathbf{x}_{0}$, the first-order gradient vanishes, and to leading order the loss surface will look like $L(\mathbf{x})\approx L(\mathbf{x}_{0})+0.5(\mathbf{x}-\mathbf{x}_{0})^{t}\mathbf{H}_{0}(\mathbf{x}-\mathbf{x}_{0})$ for positive definite hessian $\mathbf{H}_{0}$. The gradient within this region is therefore $\nabla L(\mathbf{x})\approx\mathbf{H}_{0}(\mathbf{x}-\mathbf{x}_{0})\equiv\mathbf{H}_{0}(\Delta)$. Because $\mathbf{H}_{0}$ is positive definite, we conclude that $(-\Delta)^{t}\mathbf{H}_{0}(\Delta)<0$, which means that moving the system in the $\Delta$ direction is a valid gradient descent direction. ∎
@@
The prior theorem establishes that PROFIT accomplishes precisely what it seeks to do: even if the old data are no longer available, PROFIT allows us to train as if we can still compute the full loss function of the old data. As the system moves further away from the old equilibrium, we are able to restore some of the function of that equilibrium through these regularized updates. In particular, there is a case where PROFIT leads to a trivial update. Intuitively, since the model converged, moving away from the minimum creates a gradient ($\Delta$) that points back toward it. PROFIT leverages this implicit gradient to regularize the update on the new task.
Theorem 3.2.
@@
(Stable points) Suppose a model has weights $\theta$ and $\mathbf{O}^{(\text{ref})}$ maps $\theta\mapsto\theta^{\prime}$ and $\mathbf{O}^{(ref)}$ is SGD. If we are not at a critical point of $L_{\text{new}}$, PROFIT will result in zero change in the model weights $\theta$ if and only if $\hat{\nabla}_{\theta}L_{\text{new}}=\hat{\nabla}_{\theta^{\prime}}L_{\text{new}}$, where $\hat{\nabla}$ refers to the unit vector corresponding to $\nabla$.
Proof.
@@
In this case $\Delta$ would point in the direction $\nabla_{\theta}L_{\text{new}}$, at which point $\nabla_{\theta}L_{\text{new}}\perp\Delta=0$, and the total update from PROFIT will be just a restoration to $\theta$. If the equality of the gradient does not hold, $\nabla_{\theta}L_{\text{new}}\perp\Delta\neq 0$ will lead to a non-zero total update. ∎
Corollary 3.3.
@@
(Linearity forces a stable point) In the situation defined by Theorem 3.2, PROFIT will encounter a stable point if the loss surface is perfectly linear between $\theta$ and $\theta^{\prime}$.
@@
The prior theorem and corollary demonstrate that PROFIT will fail to move the system when the loss surface becomes exactly locally linear at a point. This almost never occurs for high-dimensional loss surfaces that exist for deep models.
Theorem 3.4.
(Convergence) For a model trained with SGD $\mathbf{O}^{(\text{ref})}$, any standard optimizer for $\mathbf{O}$, and loss $L$ under PROFIT, the model is guaranteed to converge to either (1) a stable point of the system as defined by Theorem 3.2, or (2) a convergence point of $\mathbf{O}$.
Proof.
If the system does not converge to a stable point as defined in Theorem 3.2, then it will necessarily produce a valid gradient descent direction because it is an orthogonal projection of a valid gradient direction $\nabla_{\theta}L$. As such, it inherits the same stable points as $\mathbf{O}$. ∎
In summary, PROFIT enjoys the following theoretical guarantees:
  1. 1.
    By Theorem 3.1, it mitigates catastrophic forgetting relative to prior methods.
  2. 2.
    By Theorem 3.2, it introduces only rare failure modes, which require all higher-order gradients to vanish.
  3. 3.
    By Theorem 3.4, it inherits the stable points of its parent optimizers, allowing the use of state-of-the-art and future optimizers.
3.3 Hyperparameter Discussion
PROFIT introduces three main hyperparameters for fine-tuning: $n_{\text{ref}}$, $\mathbf{O}^{\text{(ref)}}$, and $\lambda_{\text{ref}}$. $n_{\text{ref}}$ controls the degree of exploration of the reference optimizer, while $\lambda_{\text{ref}}$ controls the step-size at each iteration of the reference step. The choice of $\mathbf{O}$ can be set by the practitioner according to their needs, but $\mathbf{O}^{(ref)}$ should be set to standard SGD, as the gradient calculations that drive PROFIT are the cleanest when the reference updates are simple.
A primary concern is the cost of setting $n_{\text{ref}}$, as it requires $n_{\text{ref}}$ additional optimization steps per training step. In practice, the practitioner is encouraged to start with $n_{\text{ref}}=1$, which works well in practice, and to only increase $n_{\text{ref}}$ if performance is lacking. However, we note two reasons why the potentially increased compute of PROFIT is not a large issue: (1) we find that PROFIT generally converges faster, possibly due to the positive regularization effects of the method, and (2) fine-tuning is generally run for much fewer steps than training from scratch, so the effects of this additional training time are often still very tractable. For all experiments, we use $n_{\text{ref}}=1$.
Mathematically, $\lambda_{\text{ref}}$ encodes how quickly we travel from equilibrium to study the general shape of the loss surface, so that we can make an informed decision on where to go next. In practice, we find that there is often a sweet spot somewhere between $\lambda_{main}/10$ and $\lambda_{main}/10000$, but the exact value is highly dependent on the loss surface for the specific problem. In fact, it may turn out that the optimal value of $\lambda_{\text{ref}}$ tells us something about the fundamental properties of the loss surface of that particular setting, but such analysis falls outside the scope of the present work.
3.4 Memory Considerations
One of the primary difficulties in implementing PROFIT is that storing the reference weights $\theta_{\text{ref}}$ as defined in Algorithm 1 will increase the memory footprint. In memory profiling experiments (Appendix D.1), we found that a vanilla implementation of PROFIT will on average increase the memory overhead of model training by approximately 25%, which then decreases to 0% if we remove the additional memory consumed by $\theta_{\text{ref}}$. In many fine-tuning scenarios, this additional memory consumption may not pose a problem, as model backbones may be largely frozen during fine-tuning. However, it is still worthwhile to address this substantial memory footprint.
[…]
3.5 Assumptions
@@
Orthogonalizing between $\mathbf{g}$ and $\Delta$ implies that we are interested in resolving the conflict between the original and fine-tuning datasets. If the original and fine-tuning distributions are significantly different, the orthogonalized gradient direction may lead to destructive interference. Therefore, our interpretation holds only if the original and fine-tuning datasets are from proximal distributions. Assumption 1 may differ from typical fine-tuning literature and appear greatly limiting, but is exceedingly common in practice. For example, a self-driving car may train a classification head for a new scooter type while maintaining performance on existing vehicles. Assumption 2 is required since our method is exclusively designed for fine-tuning (which is a guiding principle). Nonetheless, we include it here to emphasize to the reader that PROFIT cannot be used for general model optimization.
@@ 4 Experiments
We now detail a number of experiments for PROFIT in diverse settings: image classification, visual task adaptation, and large-scale motion prediction. We primarily focus on comparisons to standard fine-tuning with commonly used optimizers (on either the full model or just the model head), as those are - by a large margin - still the most commonly used fine-tuning methods in the industry due to their known performance and ease of implementation. We will show that PROFIT, while easy to implement, provides a significant performance boost in all cases.
@@ 4.1 A Simple Toy Example
We first apply PROFIT to a simple, 2D regression problem with MLPs as a toy example. We choose the 2D function $f(\mathbf{x})=\sin(10|\mathbf{x}|)$, as radially symmetric and periodic functions pose some challenge for neural networks models. We also add $\mathcal{N}(0,1)$ noise to the output. Even though we want to fit to a low-dimensional example, it is still important for the problem to be difficult enough to see interesting behavior within our models, especially given the dimensionality requirements as discussed in Section 3.2 of the main paper. The original dataset consists of input-output pairs where the input coordinates are drawn independently from $\mathcal{U}[-1.0,1.0]$, while the new dataset we wish to fine-tune on has input coordinates drawn from $\mathcal{U}[0.8,1.5]$. This is clearly also challenging because, although the domains of the two datasets overlap in the interval $[0.8,1.0]$, they are largely non-overlapping.
Our MLP consists of three layers with weights of shape $2,500]\rightarrow[500,500]\rightarrow[500,1]$, and we use RMSProp as our baseline optimizer. The baseline model is trained on the original data split only for 10000 steps at a learning rate of 1e21e-2, with fine-tuning runs trained at a learning rate of 5e45e-4 for 1500 steps. For exact details on the training procedure, please refer to the Appendix C.1.
Method Original Data Error $(\downarrow)$ New Data Error $(\downarrow)$
@@
Table 1: Results on a 2D toy problem. A baseline is trained on just the original data domain and then fine-tuned at a lower learning rate on a new data domain (both the full model and the head only). This is compared to PROFIT, which is trained on the full model. Fine-tuning on the model head weights only provides some protection against performance regression on the original data, but also is less able to adapt to the new data. PROFIT outperforms both baselines and shows impressive resilience in maintaining performance on the original split. All results have standard error $\leq 0.01$.
The results are visualized in the bottom half of Figure 1, with more granular visualizations in Appendix C.1. Visually, the benefits of PROFIT are pronounced; standard fine-tuning creates a warped shape bearing minimal resemblance to the original ground truth. PROFIT effectively remembers the original shape with only minimal regressions along the steep edges of the distribution.